OpenStudy (anonymous):

simplified form of 9x^2-25 / 3x-5 ?

OpenStudy (anonymous):

would the answer be 3x + 5, with the restriction x ≠ 5/ 3

OpenStudy (mathstudent55):

\(\dfrac{9x^2 - 25}{3x - 5} \) Can you factor the numerator?

OpenStudy (anonymous):

wow my answer can't beat that

OpenStudy (mathstudent55):

\(\dfrac{9x^2 - 25}{3x - 5} \) \(= \dfrac{(3x + 5)(3x - 5) }{3x - 5} \) \(= 3x + 5; ~~3x - 5 \ne 0 \) \(3x - 5 = 0 \) \(3x = 5\) \(x = \dfrac{5}{3} \) \(\dfrac{9x^2 - 25}{3x - 5} = 3x + 5; ~~x \ne \dfrac{5}{3} \) You are absolutely correct. Good job!